OpenStudy (anonymous):

I have been trying to figure out how to do this problem for a long time. What's the answer to this problem? Solve for x: -2(x + 3) = -2(x + 1) - 4. 1)2 2)3 3)All real numbers 4)No solution

OpenStudy (anonymous):

no soultion ?

OpenStudy (anonymous):

Its 3

OpenStudy (anonymous):

all real numbers because they both symplify to -2x-6

OpenStudy (anonymous):

But you can still draw it to one side

OpenStudy (anonymous):

they are equivalent

OpenStudy (anonymous):

you would get 0

OpenStudy (anonymous):

If you add 2x to one side and add the six as well it ends in x=3

OpenStudy (anonymous):

no -2x-6=-2x-6 +2x +2x -6=-6 +6 +6 0=0

OpenStudy (anonymous):

so is the answer 3 or no solution?

OpenStudy (anonymous):

real numbers

OpenStudy (anonymous):

is the awnser

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

no soultion

OpenStudy (anonymous):

all real numbers

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

because x becomes cancel...